PDP-8/L Restoration

2/26/11

We pulled the G785 module out of the backplane. This disconnects the power supply from the backplane. We connected a Variac between the AC power source and the 8/L and slowly, over the course of hours, increased the AC voltage to the 8/L power supply. The voltages looked OK, so we put the G785 back in the backplane and powered it on at full AC voltage. After a few minutes smoke poured from the rear transformer in the power supply. Another 8/L owner reported the same smoke and found a shorted power diode in his power supply.

We have another 8/L system that is in poor condition because rodents were living in it. The power supply looks OK, so we can try the power supply in the complete system next week.

7/30/11

Warren and Mike finally had a chance to work on this machine again. We pulled the smoked power supply from the incomplete PDP-8/L and installed in this system. Replacing the power supply is quite a project and involves removing the front panel and a lot of boards.

We noticed that the power supply from the donor system only had one transformer and the failed power supply had two transformers and a toggle swith instead of the keyswitch for poer on. The power supply schematic only shows one transformer, so we are speculating that the failed power supply was substantially modified during its 40 year life.

When we powered on the reassembled system we were rewarded with lights and no smoke. There are some signs of life! After a little experimenting we found that there were some serious problems with the registers. You can see the light pattern after we turned all of the switches off and pushed the Load Address switch. All of the lights should have been off, but many were on. Each of the M220 modules contains 2 bits of the registers. We swapped pairs of the M220 modules around and observed the changes in the light patterns. It looks like only two of the six M220 modules are fully functional. From some quick Googling it looks like the SN7474N chips are a common failure mode for these modules. At least those chips are common and inexpensive.

8/12/11

Ulrich N Fierz of Zurich Switzerland donated two yellow and two orange replacement keys for the front panel.

Mike bought an HP 10526T Logic Pulser, HP 10525T Logic Probe, HP 10528A Logic Clip, and an HP 10529A Logic Comparator. These are the 35 year old tools that are recommended to debug the PDP-8.

Dan found the double-sided Flip-Chip extenders so we can debug the M220 modules in the system.

8/20/11

Mike replaced the missing EXAM switch handle with one donated by Ulrich N Fierz. It works great now. We also found that one of the pivots on the INST FIELD switch is broken so we need to replace that one too with another spare from Mr. Fierz.

Warren and Mike continue debugging the M220 Major Register modules in this system. We started two weeks ago by moving the six M220 modules around in the system and observing the changes in the light patterns. We marked which modules had odd or even or both bit issues. The goal was to get the Memory Address register working because there is no way to Deposit or Examine memory without it.

Some of the Memory Address bits were always on. By observing the lights we could see that some of the Accumulator bits were always on and the corresponding bits in the Memory Address register were always on. The HP 10528A Logic Clip and the HP 10525T Logic Probe were really helpful determining that the AC Enable signal was always on. That was mixing the contents of the Accumulator with the contents of the Switch Register. After chasing the stuck AC Enable signal through the M617 in slot A9 to the M160 module in slot D9 we determined that either the DEC7450 E3 or the DEC7460 chip E4 was defective. There is no way determine which one was defective and we didn't have spare ICs so we swapped the M160 module with one from the parts PDP-8/L. After swapping the M160 the AC Enable signal was working OK. We tested the other two spare M160 modules and those worked OK too. Nice to have some spares! With the AC Enable signal behaving we determined that only one M220 had a problem with the Memory Address register. After a little debugging we found that half of the DEC7474N E4 was defective. We replaced the chip with a "new" TI SN7474N that had a date code only two years newer then the PDP-8/L. The pen in the picture at the left is pointing to the replaced SN7474. After the M220 was repaired we determined that all of the bits in the Memory Address register will load from the Switch register correctly.

After some more experimenting we found that repeated pushes of the newly installed Examine switch do not increment the Memory Address register. If you turn bit 1 of the Switch register on the part of the Memory Buffer is increment. It looks like there is something else wrong with the control logic for the Major Register boards. That will be the project for next week.

We need to get the registers working correctly before we can address any of the other problems with this system.

8/27/11

Old problems:

1. MA does not increment when SR=0, LA, EXAM.

2. MB lower bits seem to increment when SR=1; LA, EXAM,EXAM,EXAM...

2b. Two MB bits (bits 4,5?) seem to toggle together

i.e. not increment like the lower bits (0 to 3?)

Debugging problem 2:

Looking at MB11 (AB02 M220)

(Schematic D-BS-8L-0-9 Major Register Gating, Sheet 3, Zone 1 & 2)

REG BUS 11 is showing incremented value (wrong).

ADD output is also wrong.

Looking at AND-OR terms, SR ENABLE is ON (wrong) when EXAM,EXAM...

Looking at SR ENABLE

(Schematic D-BS-8L-0-4 Reg. Output Gate Control, Zone B4)

A08 M617 pin E1 is SR ENABLE.

A08 M617 pins A1,B1 (inputs) are wrong.

Driven by C12 M115 pin U1.

C12 M115 inputs P1,R1,S1 are okay.

So, C12 M115 E2 7410 is bad. Swapped in a spare. Works (does not increment).

Note: At some point, we determined SR=2 causes MB10 and up to increment.

Now, it makes sense. Since SR ENABLE was incorrectly being enabled, it was ADDing the SR to MEM read value, which was being written back to CORE (remember: CORE read is destructive, so it writes back the MB value).

So, CORE seems to read and write but always at 0000 (Problem 1: MA does not increment)

Problem 2B has changed. The affected bits seem to be stuck on.

We tried swapping:

Major register bits (AB02 thru AB07 M220s),

Sense Amps (A18,A19,A20,B18,B19,B20 G020s).

Inhibit Drivers (A23,B23,B24 G228s)

Much head scratching...(what's really in memory????)

Eventually, we decided all M220s seem okay, but 6 bits are stuck either high or low.

Current problems:

1. MA does not increment when SR=0, LA, EXAM, EXAM...

2. 6 memory bits are stuck high or low.

New problems noticed:

3. RUN -> RUN light. STOP does not stop, but SINGLE STEP does stop.

4. ION is on (related to problem 1?).

9/10/11

If you do a Load Address and then RUN, the STOP key doesn't work. Pushing SINGLE STEP will stop the processor.

Starting in section C2 on page D-BS-8L-0-2 of the DEC-8L-HR2A-D schematics, the run flip-flop is on the M216 module in slot C06 so we started looking there.

We first looked at the POWER CLEAR-N signal on pin R1. We saw a 230 ms pulse with a 30-40 ms up slope. That is OK.

Next we looked on pin N1 for the TP3 pulses. They were there with a 1.66 usec period, about 600 KHz. That is OK.

Next we looked at the M160 module in slot D09. The E1 pin went high when the STOP key was pressed. high for stop (OK). That is OK.

We looked at the F SET signal on pin F1 and found that it was always low.

We then moved to the M111 module in slot C07. It is in section D5 on page D-BS-8L-0-3 of the schematics. We found that the F SET-N signal on pin M2 was always high.

We moved to the M117 module in slot C11. It is in section C5 on page D-BS-8L-0-3 of the schematics.

The D SET-N signal on pin D1 was always high. That is OK.

The E SET-N signal on pin C1 was always high. That seems OK.

The BREAK OK-N signal in pin B1 was always high. That is OK.

The SPECIAL CYCLE-N signal on pin A1 was always low. That is wrong.

SPECIAL CYCLE is WORD COUNT or CURRENT ADDRESS. Sounds like it is part of the Data-Break facility that is not installed.

We looked at the M113 module in slot D12. It is in section B2 on page D-BS-8L-0-3 of the schematics.

The SPECIAL CYCLE signal on pin N1 was always high which is wrong. We actually found that the signal was at 1.7 V, neither high or low.

That means that the DEC7400N E2 is bad. We tagged the module for future repair.

We replaced the defective M113 with a NOS one from a recent donation. Now the SPECIAL CYCLE signal on pin N1 is always a solid low.

We traced the SPECIAL CYCLE signal back through pin D2 on the M111 module in slot D13 at section B2 on page D-BS-8L-0-3 of the schematics

The SPECIAL CYCLE-N signal on pin D2 was always high. Much better!

We looked at the M117 module in slot C11 at section C5 on page D-BS-8L-0-3 of the schematics.

The SPECIAL CYCLE-N on pin A1 was always high. Looks good!

The E SET-N signal on pin C1 looked strange.

We moved to the M115 module in slot C09 at section D4 on page D-BS-8L-0-3 of the schematics.

The E SET START signal on pin V1 pulses high once per power-up.

The INT OK-n signal on T2 is high.

The signals on U2, and V2 are high.

The current behavior is:

LINK is not cleared by START.

If you set the switches to 0000, then LOAD ADDRESS, set SINGLE STEP on, and push START; FETCH is on. That looks OK.

We probably got confused by the E SET signal behavior because the 5000 JMP and 7000 NOP instructions do not have an Execute cycle.

We set the switches to 0000, then LOAD ADDRESS, then DEPOSIT, and continued with the E SET signal.

We looked at the M117 module in slot C11 at section C5 on page D-BS-8L-0-3 of the schematics.

The F SET-N signal on pin E1 is HIGH because the BREAK OK-N signal on pin B1 is LOW. This is not correct.

We continued with the M117 module in slot C11 at section C2 on page D-BS-8L-0-3 of the schematics.

The BREAK OK-N signal on pin J2 is LOW. This is not correct.

The BRK SYNC signal on E2 is always HIGH. This is wrong.

We moved to the M216 module in slot B11 at section B2 on page D-BS-8L-0-3 of the schematics.

The BRK SYNC signal on pin V1 was always high (never pulsed low).

The MANUAL PRESET-N signal on pin U2 (clear?) pulsed low at start.

The BRK RQST-N signal on pin T2 was always high (ok).

Pin V1 is actually Q-N output and pin U2 is PRSET-N.

This means that the DEC7474N E2 is bad.We tagged the module for future repair.

We replaced the M216 module at B11 with a NOS (not sealed bag) module.

Now STOP works and the FETCH and EXECUTE lights cycle. This is some real progress.

At this point it looks like we fixed a major problem with the BRK SYNC-N signal so that the FETCH and EXECUTE cycles work, but we broke some other behavior with LOAD ADDRESS. Sometimes when we press the LA key the SW register contents are not loaded into the MA register. We will need to back out these module swaps to see if we can determine what we fixed and what we broke. I will fix the broken modules so we can put them back in the machine to see if it behaves better. We assumed, but really don't know if the NOS modules was OK.

10/08/11

During the last debug session we replaced the broken M113 in slot D12. This fixed the floating SPECIAL CYCLE signal and allowed the FETCH/EXECUTE cycles to work, but broke the Load Address switch.

We replaced the NOS M216 in slot D12 with another one, but is made no difference in the behavior.

We put the NOS M113 back in slot D12 and now LA switch works. Well, at least it worked once. Subsequent operation of the LA switch did not work correctly.

We set the switches to 0000, operated the LA, DEP, DEP 0000 LA SS

START fetches, CONT executes CONT no lights

Our guess is that fixing the SPECIAL CYCLE signal unmasked yet another problem in the 8/L.

We put the broken M113 slot D12 so we could debug LINK not cleared by START.

Looking at LINK not cleared by START.

We tried tried storing CMA CML instructions in core and executing them.

0000 LA 7020 (CMA CML) DEP MB is 7020

0000 LA EXAM doesn't read

trying to DEP at non 00, it appears that

EXAM clears MA and DEP clears MA

We put the NOS M113 back in slot D12 and verified that LA is not working.

Random MA, START does FETCH/EXECUTE.

On the M617 in slot B08 pin M1 is high, pulse low on LA (ok). all switches pulse it except CONT (ok).

Set the switches to 0000, did a LA, MB = 4010.

On the M220 in slot AB02 looking at enable signals.